solve the system of equations by graphing. Then classify the system
x+y=9
x-y=3
What is the soltion of the system. Select the correct choice below and if neccessary fill int he answer box to complete your choice.
A- The solutin is___
B- There are infintely many solutins
C-There is no soluton

Answer by mananth(16946)

x+y=9
x=0,y=9 (0,9)
y=0,x=9 (9,0)
...
x-y=3
x=0,y=-3 (0,-3)
y=0,x=3 (3,0)


The solution is (6,3)